OpenStudy (dmezzullo):

Evaluate: 4 - 49 ÷ 7 + 32 -12 2.6 3 6

OpenStudy (dmezzullo):

is it -12?

OpenStudy (anonymous):

include the brackets please

OpenStudy (dmezzullo):

there are nno brackets

OpenStudy (dmezzullo):

thats the question can u help plz

OpenStudy (anonymous):

i m trying to solve this but it is very confusing without brackets

OpenStudy (dmezzullo):

oh ok

OpenStudy (anonymous):

Answer is 29

OpenStudy (anonymous):

49 divided by 7 =7

OpenStudy (dmezzullo):

29 aint an option

OpenStudy (anonymous):

@dianarock336 i agree with you

OpenStudy (anonymous):

4-7=-3 +32=29

OpenStudy (anonymous):

Crap, are you sure!?

OpenStudy (dmezzullo):

ya thats the problem

OpenStudy (anonymous):

-45/39 can also be the answer

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

It could 3

OpenStudy (anonymous):

= 4-1 =3

OpenStudy (anonymous):

what about 2.6

OpenStudy (anonymous):

Yeah

OpenStudy (dmezzullo):

thats the answer

OpenStudy (anonymous):

Its 2.6

OpenStudy (anonymous):

are you sure thats the answer

OpenStudy (anonymous):

if it is then i can show you how it comes